OpenStudy (anonymous):

what is the solution of the equation? 3^5 sqrt(x+2)^3 +3 =27

OpenStudy (anonymous):

\[3^5 \times \sqrt{(x+2)^3} +3 =27\]What's the equation? use ( )

OpenStudy (anonymous):

That's the whole equation. That's how it's typed on my test

OpenStudy (anonymous):

i guess it wants you to figure out x

OpenStudy (anonymous):

I'm sorry I don't want to do this problem b/c it is written so poorly. It isn't clear what the equation is. Also, we are not allowed to help with tests/quizzes